OpenStudy (anonymous):

What is the missing number? –3.4 + (–8 + 2) = [ + (–8)] + 2

OpenStudy (anonymous):

what do you think it is?

OpenStudy (anonymous):

-3.4

OpenStudy (anonymous):

you got it

OpenStudy (anonymous):

help more [lz

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Which property could be used to find the missing number? –4 • (–6.2 + 5) = –4 • + (–4) • 5 A. associative property B. distributive property C. commutative property

OpenStudy (anonymous):

but i will need to leave in about 15 minutes.

OpenStudy (anonymous):

ok this need help

OpenStudy (anonymous):

can you type in the numbers?

OpenStudy (anonymous):

what do u mean by the number

OpenStudy (anonymous):

the question is in diamonds

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

–4 • (–6.2 + 5) = –4 • + (–4) • 5

OpenStudy (anonymous):

-4*(-6.2+5)=-4*_+(-4*5)

OpenStudy (anonymous):

4.8 is the value for the first equation right?

OpenStudy (anonymous):

what do you think it is?

OpenStudy (anonymous):

i gotta go tr

OpenStudy (anonymous):



OpenStudy (anonymous):

ok back @CaseyBabe

OpenStudy (anonymous):

u there???